OpenStudy (anonymous):

Identify the center and radius for the following equation: x²+y²-6x+10y-15=0

OpenStudy (tkhunny):

I let the last one go, but now that you've done it again, I'll venture. "Equations" don't have either a center or a radius. Graphs related to equations might. Your task is to complete the square in x and in y, allowing the constant term to be what it may choose to be. x²+y²-6x+10y-15=0 x² - 6x + y² +10y = 15 x² - 6x + _____ + y² +10y + _____ = 15 + _____ + _____ Fill in the blanks!

OpenStudy (anonymous):

i'm not sure, thats what the question asks

OpenStudy (tkhunny):

Well, sorry. Whoever wrote the question should make a better effort to speak precisely. So, are you going to fill in the blanks? Let's see it.

OpenStudy (anonymous):

yeah, idk thats the question on the worksheet i got, i'm not sure what to fill in :( @tkhunny

OpenStudy (anonymous):

@erin512 Hint: use (a+b)^2=a^2+b^2+2ab